# Addition is almost all you need: Compressing neural networks with double binary factorization

### v2

@@ Abstract
Binary quantization approaches, which replace weight matrices with binary matrices and substitute costly multiplications with cheaper additions, offer a computationally efficient approach to address the increasing computational and storage requirements of Large Language Models (LLMs). However, the severe quantization constraint (±$\pm 1$) can lead to significant accuracy degradation. In this paper, we propose Double Binary Factorization (DBF), a novel method that factorizes dense weight matrices into products of two binary (sign) matrices, each accompanied by scaling vectors. DBF preserves the efficiency advantages of binary representations while achieving compression rates that are competitive with or superior to state-of-the-art methods. Specifically, in a 1-bit per weight range, DBF is better than existing binarization approaches. In a 2-bit per weight range, DBF is competitive with the best quantization methods like QuIP# and QTIP. Unlike most existing compression techniques, which offer limited compression level choices, DBF allows fine-grained control over compression ratios by adjusting the factorization’s intermediate dimension. Based on this advantage, we further introduce an algorithm for estimating non-uniform layer-wise compression ratios for DBF, based on previously developed channel pruning criteria.
@@ 1 Introduction
Figure 1: Left: Schematic drawing of computation in Double binary factorization. $X$ is the layer input in FP16 format, $W$ is the original weight matrix. $A$ and $B$ consist of ±$\pm 1$ elements, $a,m,b$ are vectors in FP16 format. We omit matrix transpositions for simplicity. Right: Comparison of our Double binary factorization (DBF) with previously proposed network compression methods on Llama2-7B.
@@ 3 Double binary factorization
where $a,b$ are column vectors,$W_{\pm 1}$ is a sign matrix with entries from $\{-1,1\}$ and $\odot$ denotes elementwise (Hadamard) product with appropriate broadcasting. During linear layer computation, we calculate:
@@ 3.3 Input and output importance scaling
While many compression schemes try to solve the layer-wise pruning problem ($\min||XW-XW_{c}||_{2}^{2}$, where $X$ is the calibration input, $W$ the original matrix, and $W_{c}$ the compressed matrix) (Frantar and Alistarh,, 2023; Frantar et al.,, 2022; Tseng et al., 2024a, ), we opt to use a slightly different approach.
@@ 3.4 Fine-tuning during and after factorization
Almost all state-of-the-art quantization algorithms use some form of fine-tuning. We use slight variation of the procedure from QuIP# (Tseng et al., 2024a, ) and QTIP (Tseng et al., 2024b, ) as follows: We use a small calibration dataset (in our case, 256 sequences). When compressing the $i$-th transformer block, we collect its expected output $Y^{(i)}$ from the original dense model. We also collect the output of the $(i-1)$-th block after compression, which becomes the input $X^{(i)}$ of the $i$-th block. Before compression, we first fine-tune the block to correct errors from previous blocks. Then we compress the query, value, and output matrices, fine-tune the rest of the block, and compress the remaining matrices. After the compression, we fine-tune the remaining unquantized parameters (scaling vectors in each binary factorization).
@@ 4.1 Uniform compression of LLMs
Table 1: Results for uniform compression of Llama2-7B. Avg. bits refer to the average number of bits per original weight.
@@
Results. Results for uniform compression are shown in Tab. 1, and 2. Our results are comparable to AQLM with PV-tuning for 2.3 bit compression. We are marginally worse than QTIP but better than QuIP# for 2 bit compression. Both methods require decompression into full precision weight, while DSF uses mostly additions. For 1-1.5 bit compression, our method is superior to other tested methods. Note that OneBit was fine-tuned for much longer than DBF (OneBit reports 7 days on 8 H100s, we tune for 1 day on 4 A100s), and DBF with only continuous parameters fine-tuning is still better. Also, note that to achieve the average 1 bit per layer, we need the middle dimension of factorization to be smaller than the dimensions of the original matrix. But even with the low-rank bottleneck, DBF is still better than OneBit.
@@ 4.2 Non-uniform compression of LLMs
Table 3: Comparison of uniform and nonuniform compression of Llama3-8B.
Avg. bits Method Wikitext ppl ArcC ArcE PiQA Wino Avg. zero shot
16 Dense 5.54 50.43 80.09 79.71 72.61 70.71
2 QTIP 7.33 44.20 75.20 77.60 70.70 66.93
2 DBF + PV 7.39 41.38 74.11 77.04 67.40 64.98
2 DBF + PV + NU 7.39 41.80 74.32 77.36 68.19 65.42
Due to a lack of computational resources, we run only one round of iterative pruning. We compress Llama3-8B in the same way as above. We start with DBF with 2.1 bits/weight. We then estimate scores for middle elements of the factorization and compute sizes of each layer. We found that restricting each layer to have at least 1.5 bits/weight leads to slightly better results. We then repeat the compression and fine-tuning process from the start. Results are reported in Tab. 3. We see that a nonuniform distribution gives benefit even with one iteration of redistribution.
@@ Acknowledgments
This research was supported by grants 1/0140/25, and 1/0538/22 from Slovak research grant agency VEGA. Part of the research results was obtained using the computational resources procured in the national project National competence centre for high performance computing (project code: 311070AKF2) funded by European Regional Development Fund, EU Structural Funds Informatization of society, Operational Program Integrated Infrastructure. We acknowledge the EuroHPC Joint Undertaking for awarding this project access to the EuroHPC supercomputer LEONARDO, hosted by CINECA (Italy) and the LEONARDO consortium through an EuroHPC Benchmark and AI Access calls.

### v3

@@ Abstract
Binary quantization approaches, which replace weight matrices with binary matrices and substitute costly multiplications with cheaper additions, offer a computationally efficient approach to address the increasing computational and storage requirements of Large Language Models (LLMs). However, the severe quantization constraint ($\pm 1$) can lead to significant accuracy degradation. In this paper, we propose Double Binary Factorization (DBF), a novel method that factorizes dense weight matrices into products of two binary (sign) matrices, each accompanied by scaling vectors. DBF preserves the efficiency advantages of binary representations while achieving compression rates that are competitive with or superior to state-of-the-art methods. Specifically, in a 1-bit per weight range, DBF is better than existing binarization approaches. In a 2-bit per weight range, DBF is competitive with the best quantization methods like QuIP# and QTIP. Unlike most existing compression techniques, which offer limited compression level choices, DBF allows fine-grained control over compression ratios by adjusting the factorization’s intermediate dimension. Based on this advantage, we further introduce an algorithm for estimating non-uniform layer-wise compression ratios for DBF, based on previously developed channel pruning criteria.
@@ 1 Introduction
Figure 1: Left: Schematic drawing of computation in Double binary factorization. $X$ is the layer input in FP16 format, $W$ is the original weight matrix. AA and $B$ consist of $\pm 1$ elements, $a,m,b$ are vectors in FP16 format. We omit matrix transpositions for simplicity. Right: Comparison of our Double binary factorization (DBF) with previously proposed network compression methods on Llama2-7B.
@@ 3 Double binary factorization
[…]
@@ 3.3 Input and output importance scaling
[…]
@@ 3.4 Fine-tuning during and after factorization
Almost all state-of-the-art quantization algorithms use some form of fine-tuning. We use slight variation of the procedure from QuIP# (Tseng et al., 2024a, ) and QTIP (Tseng et al., 2024b, ) as follows: We use a small calibration dataset (in our case, 256 sequences). When compressing the $i$-th transformer block, we collect its expected output $Y^{(i)}$ from the original dense model. We also collect the output of the (i1)(i-1)-th block after compression, which becomes the input $X^{(i)}$ of the $i$-th block. Before compression, we first fine-tune the block to correct errors from previous blocks. Then we compress the query, value, and output matrices, fine-tune the rest of the block, and compress the remaining matrices. After the compression, we fine-tune the remaining unquantized parameters (scaling vectors in each binary factorization).
@@ 4.1 Uniform compression of LLMs
Table 1: Results for uniform compression of Llama2-7B. Avg. bits refers to the average number of bits per original weight.
@@
Results. Results for uniform compression are shown in Tab. 1, and 2. Our results are comparable to AQLM with PV-tuning for 2.3 bit compression. We are marginally worse than QTIP but better than QuIP# for 2 bit compression. Both methods require decompression into full precision weight, while DBF uses mostly additions. For 1-1.5 bit compression, our method is superior to other tested methods. Note that OneBit was fine-tuned for much longer than DBF (OneBit reports 7 days on 8 H100s, we tune for 1 day on 4 A100s), and DBF with only continuous parameters fine-tuning is still better. Also, note that to achieve the average 1 bit per layer, we need the middle dimension of factorization to be smaller than the dimensions of the original matrix. But even with the low-rank bottleneck, DBF is still better than OneBit.
@@ 4.2 Non-uniform compression of LLMs
Due to a lack of computational resources, we run only one round of iterative pruning. We compress Llama3-8B in the same way as above. We start with DBF with 2.1 bits/weight. We then estimate scores for middle elements of the factorization and compute sizes of each layer. We found that restricting each layer to have at least 1.5 bits/weight leads to slightly better results. We then repeat the compression and fine-tuning process from the start. We are able to push perplexity for Llama3-8B further down from 7.30 to 7.26. We see that a nonuniform distribution gives benefit even with one iteration of redistribution.
@@ Acknowledgments
This research was supported by funding from the Slovak Research and Development Agency grant APVV-24-0045 and by grants 1/0140/25, and 1/0538/22 from the Slovak research grant agency VEGA. Part of the research results was obtained using the computational resources procured in the national project National competence centre for high performance computing (project code: 311070AKF2) funded by European Regional Development Fund, EU Structural Funds Informatization of society, Operational Program Integrated Infrastructure. We acknowledge the EuroHPC Joint Undertaking for awarding this project access to the EuroHPC supercomputer LEONARDO, hosted by CINECA (Italy) and the LEONARDO consortium through an EuroHPC Benchmark and AI Access calls.